DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, April 19 Padelford C-36 3:45 pm

Singular sets of harmonic functions

Qing Han (University of Notre Dame)

We will discuss the geometry of the singular sets of harmonic functions. We will show the explicit measure estimates of these sets in terms of frequency functions. The method is from several complex analysis. It can be generalized to analytic solutions to elliptic equations.